OpenStudy (anonymous):

what is the excluded value? y^2-y+5/y+4

OpenStudy (anonymous):

If you are asking the excluded value in the domain, check for what value of 'y' the expression is meaningless.

OpenStudy (anonymous):

how do I find the meaningless "y"

OpenStudy (anonymous):

put y= -4 and check..

OpenStudy (phi):

there is a rule in math: no divide by 0. If you try to divide by 0 you get an undefined result. so if you are ever given a problem with division, you make a point of noting when or if the denominator can be 0. in your case, \[ \frac{y^2-y+5}{y+4} \] notice that if y= -4, you get a 0 in the bottom. No divide by zero rule means, *exclude y= -4* (a fancy way of saying, "y is not allowed to be -4... if it is, you can't use this expression"

OpenStudy (phi):

I assume you can figure out that if you have y+4 and y is -4 you get -4+4 and -4+4 is 0